OpenStudy (anonymous):

A camera and case together cost $110. If the camera cost $100 more than the case, how much does the case cost? Explain your answer and the steps taken to get the answer. Give your best constructed response answer.

OpenStudy (anonymous):

10 dollars.

OpenStudy (studybird123):

Well then the case would cost 10 dollars

OpenStudy (studybird123):

because the camera is 100 dollars more

OpenStudy (anonymous):

it's just 110-100. lol

OpenStudy (studybird123):

so 110-100

OpenStudy (studybird123):

= 10

OpenStudy (studybird123):

get it

OpenStudy (studybird123):

how come i dont get a medal and i answered your question whatever!

OpenStudy (nasis):

Im not sure, but to have a difference of 100, x=camera y=case x=100+y x-y=100, there is many ways to solve this, but lets keep it simple, if it was, x=100 and y =10 then x-y=90, so its false, now if x=105 and y=5, then x+y=110 and x-y=100. So the case costs 5 and the camera 105. I might sound stupid, but hey, im just a junior high school student from Greece xD.